CalCOFI 1501NH winter cruise on RV New Horizon; 15 Jan - 8 Feb 2015. 100 stations were successfully occupied; sta 60.90 & a few net tows were cancelled due to high winds & rough seas. SCCOOS stations 80.0 50.5 & 85.4 35.8 were skipped. Refer to station activities for specific station work performed.
CTD notes: due to the inability of the CTD to power dual T, C, & O2 sensors after sta 64 (80.90), single T, C, & O2 data were collected on casts 65 -100. These include sta 80.100 and all stations north on Lines 77 - 60.
After the departure of 8 scientists & volunteers in Monterey, the CTD-rosette was programmed to trip 12 bottles. Refer to the CTD Summary Notes when they are available.

Data Files:   1501NH Preliminary Plots; 1501NH CTD (raw) Cast files; 1501NH FinalQC Bottle-corrected CTD Files
Ancillary Data Links:  NAVOCEANO Drifter Data Maps
Announcement:  CalCOFI 1501NH Cruise Announcement PDF
Cruise Dates: Thursday, 15 Jan 2015 - Sunday, 08 Feb 2015
Loading: Monday, 12 Jan 2015 - Wed, 14 Jan 2015 ; 10th Ave Marine Terminal, San Diego, CA
Offloading:  09 Feb 2015; 10th Ave Marine Terminal, San Diego, CA
Ship: RV New Horizon
Station plan: 109 Stations (100 CalCOFI + 9 SCCOOS)
Cruise Map:    
Contacts: Dave Griffith, Chief Scientist, NMFS
  Amy Hays, NMFS
  Jennifer Rodgers-Wolgast, IOD (SIO Technical Coordinator)

Please note that the anomalies are based on 50-year harmonics (1949-1999) and these data are 1m binavg Seasoft-processed CTD data that have not gone through thorough quality control checks or bottle corrections. They are presented to give a general impression of current conditions. Final data plots will be published after all data are processed.


CTD Processing Summary CalCOFI 1501NH CTD Final Data
Download 1501NH CTD raw cast files zipped   Download 1501NH FinalQC CTD + bottle data
Note: Metadata files compiled from CTD processing are split 1-64 (dual TCO), 65-100 (single TCO)
General CTD Notes - data acquisition notes, logistics, processing - see below
CTD sensor corrections derived by comparing 4 secs of CTD sensor data (prior to bottle closure) to bottle samples
Dual T, S, & O2 Primary Sensor Secondary Sensor
Salinity offset (bottle - CTD salinity; > 350m only; Seabird SBE4) -0.00232 0.00084
Oxygen ml/L (dual Seabird SBE43) y = 1.0507x - 0.0094
R² = 0.9997
y = 1.0397x - 0.0132
R² = 0.9998
Oxygen umoles/Kg (dual Seabird SBE43) y = 1.0507x - 0.4099
R² = 0.9997
y = 1.0397x - 0.5732
R² = 0.9998
Single sensors Linear Polynomial
Nitrate - ISUS 4sec ave voltage vs Bottle Nitrate (Satlantic MBARI-ISUS v2) y = 27.23x - 7.1464
R² = 0.9945
Fluorometer - linear & polynomial regressions

y = 7.1988x - 0.2187

R² = 0.8363

y = 7.2289x2+5.7762x-0.1677

R² = 0.839

  Secondary CTD O2 Sensor vs Bottle O2
MBARI-ISUS Voltage vs Bottle NO3 Wetlabs ECO/FL Fluorometer Voltage vs Chlorophyll-a
General notes: These are cast & final CTD Processing Notes from 1501NH cruise

CalCOFI 1501NH General Cast Notes: This cruise occupied 100 stations: 93 standard, & 7 SCCOOS.  The stations were occupied in traditional sequence but MBARI-SECRET 20nm stations, 63.65 63.75 63.85 & 63.95 were not occupied. SCCOOS stations 80.0 50.5 (too rough, windy night) & 85.4 35.8 (offline, shiptime) were not occupied. No CTD casts deeper than 770m were performed - no 1000m MBARI-SECRET casts on Line 63 & no 3500m casts on 90.90 or 80.90.
CalCOFI 1501NH was on SIO RV New Horizon. STS' two-wire termination was used intially but after three short but problematic casts, a new termination was done. 150m of severely rusty wire was removed then a fresh termination of the three-conductor wire done by JRW. As recommended by STS, only the white conductor and shield were used for signal & ground (as the first three stations). There were problems with the pumps failing after 50m so the secondary pump was disconnected. For casts 04-11, only the primary sensors T, C, & O2 were collecting good data. Since using a single conductor for signal proved unreliable, the black+white conductors were combined for signal, shield as ground and new secondary pump installed. This solution worked and both sets of T C O2 sensors logged data for casts 12 - 63. On cast 64, the CTD failed during the downcast. The only viable workaround was to run primary T, C, O2 & pump only for casts 65 - 100. It is uncertain whether the CTD fish or deck unit were having problems or the conductive wire. (Note: CTD fish and deck unit tested fine so problems are attributed to the wire quality). The CTD & deck unit were sent to Seabird post-cruise for service. Please note that wildedit was applied to the CTD data to despike the sensor data - there were a few casts with significant spikes resulting in data gaps in 1m averaged data. Downcasts should be good but there are a few upcasts with spikes when triggering bottle closure. Please refer to specific cast notes below.

CTD initial configuration was standard: Seabird 911+ with dual T, C, O2, & pumps;  Wetlabs C-Star 25cm transmissometer; Biospherical QSP200L PAR;  Datasonics/Benthos Altimeter; WET Labs ECO-AFL/FL; Seabird  pH; Satlantic ISUS v2 & battery. Please refer to the xmlcon files or cruise  prospectus for additional info. The secondary pump was disabled for a couple casts at the beginning of the cruise; the secondary set of sensors were disabled after cast 64 for the remainder of the cruise.

Cast Notes:

Cast 01 - non-confirmations for bottles #2 at 50m and #3 at 42m; suspect #1 was 60m then #2 was 36m, both prodo bottles. Noisy signal & spikes noted on upcast; downcast looks okay.

Cast 02 - SCCOOS shallow station immediately after Cast 01; signal noise but 4 bottles marked, tripped, & closed.

Cast 03 - SCCOOS shallow station; downcast prfiles looks okay; upcast: 5m bottle #4 did not close - carousel jumped from 4 to 2 so additional trips were necessary to catchup. CTD reterminated post-cast.

Cast 04 - 93.28 severe problems, lost 12hrs troubleshooting CTD issues; CTD reterminated after 150m of wire removed. CTD pumps are turning off at ~50m on downcast. Additional solutions attempted: 1) 1° conductivity sensor changed before cast; 2) carousel changes as per DAG suggestion; 3) new pump cable; 4) new pylon cable; 5) spare deck unit used; 6) dummied secondary pump WORKED. So secondary sensor data on casts 004 - 011 are bad.

Cast 04 - 11  primary sensor data good, secondary sensor bad - no pump

Cast 06 - bottom valves on bottle #1 & #20 replaced post-cast, drippy

Cast 08 - bottle #10 replaced, bottom o-ring leaking

Cast 12 - backup pump installed as 2°; dual conductors - black & white combined for signal, shield as ground. Both pumps & sensor sets are working.

Cast 15 - rockin-rolling

Cast 19 - 15kt winds but relatively calm

Cast 20 - some big rolls

Cast 21 - rolling

Cast 22 - windy, rolling

Cast 23 - going down slow 30m/min then 40m/min; sunny but windy & rough

Cast 24 - going down slow 30 then 40 - rough seas

Cast 25 - mid-size rollers

Cast 27 - calmer; two markers at 440m - file edited/corrected

Cast 29 - calm sunny

Cast 37 - calm bright but cloudy; transmissometer dropout @ 500m downcast

Cast 40 - small pH spike at 330m on downcast

Cast 45 - calm morning; Fukushima water collected from underway since all bottle needed

Cast 48 - no nav file; created one from marker file

Cast 60 - RS232 issues, deck unit reset seemed to fix the glitch; MET system froze & was reset as well. Extra 4om tripped & ignored

Cast 61 - relatively calm with some big rolls - trough

Cast 62 - deck unit error lights flashed red at 35m downcast; spikes in O2 135m downcast; fluorometer spike @ 458m upcast. Connectors checked and serviced post-cast.

Cast 63 - all sensors glitched out at 85m on downcast but transient

Cast 64 - carousel reset after major spike, RJ did not fire enough bottles to "catch-up" so no bottles closed less than 200m ie only the 1st 7 depths from 515 to 200 trapped water. Downcast does not look bad but the transmissometer dropped out at 450-515m downcast. Pre-cast notes by RJ: deck unit displaying zeros on all channels, changed sea cable slo-blow fuse & fast-blow fuse. Numbers came back but it was decided to run only primary sensors and disable secondary sensors & pump.

Cast 65T - test cast, primary T, C, O2 & pump only on white & black conductors. Looked at sliprings but did not see any obvious problem. Test successful.

Cast 65 - single sensors no problems at all

Cast 66 - Fukushima seawater collected at 25m bottle 19 (should have tripped another for 20L; used remainder of bottle 18) and surface bottle 22. Pr offset adjusted by -0.05.

Cast 69 - rough seas again

Cast 70 -71 - moderate rollers

Cast 73 - no nav file created - will create one from marker

Cast 75 - big rollers, transmissometer failure 450-515m downcast.

Cast 76 - big rollers, hard to hit target depths; transmissometer out at 500-515m

Cast 77 - big rollers, hard to hit target depths; transmissometer out at 480-515m

Cast 78 - big rollers, hard to hit target depths; transmissometer out at 480-515m

Cast 79 - rough sunny weather; Windex bottle broke in wetlab (ammonia?); transmissometer okay

Cast 80 - moderate seas; transmissometer dropped out ~450-515m downcast

Cast 81 - Transmissometer spikes ~430-490m

Cast 82 -84 - Transmissometer squirrely ~450-500m downcast; no spare (Goericke's spare uses older connector)

Cast 85 - crossed lanyards on bottles #2-3

Cast 86 - 100 - 12 bottle casts, Leg II after 8 scientists disembarked in Santa Cruz/Monterey

Cast 88 - 100 - Transmissometer squirrely ~400-515m downcast; no spare (Goericke's spare uses older connector)

Cast 92 - moderate rollers

File notes:
08Mar2017 update to 20-1501NH_CTDFinalQC.zip file:

Seasoft-generated asc-hdr files are available renamed to 20-1501NH_LLLLSSSS_###d or u.asc & .hdr. Voltages in the .asc files were also relabled. Since this makes it difficult to reprocess & merge with bottle data if necessary, the original .asc, .hdr, & .btl have also been archived. Their voltage header labels have not been changed so refer to the corresponding .hdr file for sensor configuration. This cruise had both single T,C & O2 and dual T,C, & O2 - refer to notes.

Missing nav files created from mrk files: Cast 41, 62, & 73

Mislabels found and corrected:

Cast 09 - cast mislabeled 08 in hdr, hdr & hex files corrected using text & hex editors.
Cast 68 - sta mislabeled 90.0 in hdr, should be 80.0; hdr & hex files corrected using text & hex editors.

26Feb2015 / 16Aug2016(final update) / 08Mar2017 (CTD.csv file renaming; .asc file sensor-voltage relabeling)

Ten drifters provided by NAVOCEANO were deployed on CalCOFI 1501NH.
Real-time data may be viewed at: http://gdp.ucsd.edu/calcofi

Please note that drifters were not deployed in numerical order ie 139476 was deployed 1st then 139477 139469 139470....
Refer to date, time, CalCOFI Line & Sta for launch sites. This table is an excerpt from CalCOFI's Cruise Event Log. Click [Read More] for deployment map
and drifter track as of 17 Feb 2015.

Cruise Event# CCLine CCSta DatePST TimePST Latitude Longitude Activity Event Operator DateUTC TimeUTC Drifter#
1501 517 86.7 70 23-Jan-15 09:55:36 32.6594 -121.034 Ancillary DRIFTER DEPLOYED JRW 23-Jan-15 17:53:44 139476 - 1ST DRIFTER
1501 537 86.7 90 23-Jan-15 21:18:06 31.99437 -122.397 Ancillary DRIFTER DEPLOYED JRW 24-Jan-15 05:16:44 139477
1501 561 86.7 110 24-Jan-15 09:32:14 31.32323 -123.746 Ancillary DRIFTER DEPLOYED JLW 24-Jan-15 17:29:29 139469
1501 807 76.7 100 28-Jan-15 16:51:58 33.38861 -124.323 Ancillary DRIFTER DEPLOYED AK 29-Jan-15 00:50:29 139470
1501 816 76.7 90 28-Jan-15 23:01:23 33.7237 -123.653 Ancillary DRIFTER DEPLOYED AK 29-Jan-15 06:59:59 139472
1501 826 76.7 80 29-Jan-15 05:29:29 34.05948 -122.948 Ancillary DRIFTER DEPLOYED 139471 RAJ 29-Jan-15 13:27:59 139471
1501 843 76.7 70 29-Jan-15 12:05:51 34.3901 -122.255 Ancillary DRIFTER DEPLOYED SMM 29-Jan-15 20:04:29 139468
1501 967 70.0 90 31-Jan-15 19:36:19 34.88646 -124.479 Ancillary DRIFTER DEPLOYED SMM 01-Feb-15 03:34:59 139475
1501   60.0 80.0 05-Feb-15 11:30:00 36.39333 -125.053 Ancillar DRIFTER DEPLOYED SMM 05-Feb-15 19:30:00 139473
1501 1195 64.3 74.1 05-Feb-15 18:11:03 36.39848 -124.083 Ancillary DRIFTER DEPLOYED DAG 06-Feb-15 02:09:44 139474

Cruise Status (Twitter)

Go to top
Our website is protected by DMC Firewall!